OpenStudy (anonymous):

Another system of equation.. Need Help anyone? x+y+z=7 x+2y-z=0 x+5y=7

OpenStudy (anonymous):

Help anybody??

OpenStudy (anonymous):

x+y+z=7 (1) x+2y-z=0 (2) x+5y=7 (3) (3) gives y=1/5(7-x) (2) gives x+2/5(7-x)-z=0 3/5x+14/5-z=0 z=3/5x+14/5 (1) gives x+1/5(7-x)+3/5x+14/5=7 x=2 y=1/5(7-x) y=1 z=3/5x+14/5 z=4 The system of equations has got an unique solution. x=2 y=1 z=4

OpenStudy (nurali):

x+y+z=7----(1) x+2y-z=0---(2) x+5y=7-----(3) add '1' and '2' 2x+3y=7 2x=7-3y x=7-3y/2----(4) put 'x' in equation '3' 7-3y/2+5y=7 7-3y+10y=14 7y=7 y=1 put 'y' in equation '4' x=7-3/2 x=4/2 x=2 put x and y in equation '1' 1+2+z=7 z=4

OpenStudy (anonymous):

Thank you!!

OpenStudy (nurali):

Anytime.